OpenStudy (anonymous):

If f(x) = 7x - 1, what does f(12) represent? A.) The value of x when (7x - 1) = 12 B.)The value of (7x - 1) when x = 12 C.) The value of 12 f(x) when x = 12 D.) The value of f(x) when (7x - 1) = 12

OpenStudy (nikato):

B, f(x) where the number in the parenthesis is what u substitute for x in ur expresstion

OpenStudy (anonymous):

what you are doing in f(12)is plugging 12 in for x when 7x - 1 = 12 7x = 12 + 1 x = 13 / 7 the value of ( 7x - 1 ) when x = 12 7 ( 12 ) - 1 84 -1 83 because f(12) = translating 12 to x

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

yup